OpenStudy (ilovecake):

Six friends together earn $167.88 selling crafts. How much money will each friend get if they share the money equally? Complete the equation to model the problem. Use m to represent the unknown value. _____ = _____ 6

OpenStudy (anonymous):

\[\frac{ total\space amount }{ 6 } \] to share money equally

OpenStudy (ilovecake):

I don't get it

OpenStudy (anonymous):

total amount earned by 6 friends= $167.88 total amount earned by 1 friend = $167.88/6 m = 167.88/6

OpenStudy (ilovecake):

thanks

OpenStudy (anonymous):

anytime :)